OpenStudy (anonymous):

A rectangular garden has a perimeter of 172 feet. The length of the garden is 24 feet more than the width. What are the dimensions of the garden? (Recall the formula P = 2L + 2W.)

OpenStudy (campbell_st):

well let the width be w then the length is 24 + w so the perimeter is 172 = 2(w + 24) + 2w now you have an equation in one unknown. just distribute to remove the brackets, collect like terms then solve for w. after finding w, the length will be l = w + 24